OpenStudy (anonymous):

Simplify: (sin Θ − cos Θ)2 + (sin Θ + cos Θ)2

OpenStudy (anonymous):

@MrNood

OpenStudy (mrnood):

soz - can't help right now

OpenStudy (anonymous):

thx anyways

OpenStudy (mrnood):

the answer is simply an integer (I graphed it) but haven't got time ot prove it analytically

jhonyy9 (jhonyy9):

not sin @ = x and cos @ = y so then will get (x-y)^2 +(x+y)^2 = x^2 -2xy +y^2 +x^2 -2xy +y^2 = 2x^2 +2y^2 = 2(x^2 +y^2)= =2(sin^2 @ +cos^2 @) = 2*1 =2 (x-y)^2 = x^2 -2xy +2 (x+y)^2 = x^2 +2xy +y^2 hope this will help you sure easy

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

i dont understand

jhonyy9 (jhonyy9):

why ? this is very easy or you dont know formula sin^2 x +cos^2 x =1 ?

jhonyy9 (jhonyy9):

sorry need being y^2 (x-y)^2 = x^2 -2xy +2 the last term is y^2

OpenStudy (anonymous):

now im even more confused

OpenStudy (vishweshshrimali5):

Ok @Jessievb lets start from beginning. Ok ?

OpenStudy (anonymous):

ok

OpenStudy (vishweshshrimali5):

We have to simplify \[\large{(\sin {\theta} - \cos \theta)^2 + (\sin \theta + \cos \theta)^2}\]

OpenStudy (vishweshshrimali5):

First lets expand the square terms. We use this formula: \[\large{(a+b)^2 = a^2 + 2ab + b^2}\] Any doubt ?

OpenStudy (anonymous):

no

OpenStudy (vishweshshrimali5):

We use the above formula to expand the term - \(\large{(\sin \theta + \cos \theta)^2}\) Similarly we use the formula : \(\large{(a-b)^2 = a^2 - 2ab + b^2}\) to expand the term:\(\large{(\sin \theta - \cos \theta)^2}\)

OpenStudy (vishweshshrimali5):

Thus: \[\large{(\sin \theta - \cos \theta)^2 = \sin^2 \theta - 2\times \sin \theta \times \cos \theta + \cos^2 \theta}\] And : \[\large{(\sin \theta + \cos \theta)^2 = \sin^2 \theta + 2\times \sin \theta \times \cos \theta + \cos^2 \theta}\] Any doubt ?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

is the answer going to be a number?

OpenStudy (anonymous):

@vishweshshrimali5

OpenStudy (anonymous):

@quickstudent @WiredCardinal08 @phi @cwrw238

OpenStudy (mrnood):

@Jessievb You really do have ALL the information given above to solve this. In the time it takes for someone else to answer you - oyu could try doing it your self ignore the sin and cos for now (a+b)^2 + (a-b)^2 simplifies to 2(a^2+b^2) NOW put the sin and cos back in - and take it from there....

OpenStudy (anonymous):

ive been trying for the past half hour! so the answer is 2

OpenStudy (anonymous):

??

OpenStudy (mrnood):

Show how you get that answer...

OpenStudy (anonymous):

sin and cos = 1 i think

OpenStudy (mrnood):

??

OpenStudy (anonymous):

doesnt sin and cos cancel each other out or something

OpenStudy (mrnood):

that doesn't sound like much of a mathematical analysis! Try again???

OpenStudy (anonymous):

is the answer two? its my homework and my teacher grades hw. Ill ask her to explain it to me cause im confused